# Seeing and Knowing in the Wild: Open-domain Visual Entity Recognition with Large-scale Knowledge Graphs via Contrastive Learning

### v1

@@ Acknowledgement
The following sections present detailed descriptions of the datasets and training procedures for each experimental scenario. We outline the experimental settings for each variant of our approach, including different fusion strategies for image and text embeddings, the score functions employed by various knowledge graph embedding (KGE) methods, and key hyperparameters. Additionally, we introduce our approach for extracting efficient subgraphs from Wikidata. Finally, we provide representative examples of multi-modal prior knowledge utilized in our KnowCoL approach for a clearer overview. The pseudo code for KnowCoL algorithms are listed in Algorithm 2 and  3.
@@
OVEN dataset consists of 14 existing datasets, grounding their labels to Wikidata ids. These 14 datasets are originally created for image recognition/retrieval, and visual question answering. They are ImageNet21k-P (Ridnik et al. 2021), iNaturalist2017 (Horn et al. 2018), Cars196 (Krause et al. 2013), SUN397 (Xiao et al. 2010), Food101 (Bossard, Guillaumin, and Gool 2014), Sports100, Aircraft (Maji et al. 2013), Oxford Flower (Nilsback and Zisserman 2008), Google Landamarks v2 (Weyand et al. 2020), VQA v2 (Goyal et al. 2019), Visual7W (Zhu et al. 2016), Visual Genome (Krishna et al. 2017), OK-VQA (Marino et al. 2019), and Text-VQA (Singh et al. 2019).
@@
In this section, we elaborate the different fusion function in detail, namely addition, multi-layer perception (MLP), transformer encoder (TE), and transformer encoder with image patch embeddings (TE+IP).
@@
The default fuser we used is simple addition operation, as it shows the best generalization ability, especially for unseen entities. The fused embedding is defined as follows:
@@
The variant transformer encoder (TE) fuses the image embedding and the text embedding with transformer encoder on top. Both embeddings are fed into the encoder, allowing them to attend to each other through self-attention. The first output embedding from the encoder is used as a global representation, capturing the fused multi-modal information.
@@
Compared to the previous approach, the local image patch embeddings from the image encoder are also fed into the transformer encoder. Similarily, the first output embedding from the encoder is used as a global representation, capturing the fused information.
@@
Here, $\phi(\cdot)$ and $\psi(\cdot)$ are functions which maps nodes and relation types to corresponding embeddings.
@@
TransH is an extension for TransE, which defines relation-specific hyperplane instead of directly translation in the entity space. It can be realized by altering the score function with
@@
where $\mathbf{w}_{r}$ defines the normal vector of the hyperplane. We still use the cosine similarity and cross entropy-based approaches to conduct contrastive learning.
@@
where $\mathrm{diag}(\mathbf{r})\in\mathbb{R}^{d\times d}$ is the diagonal matrix with $\psi(r)$ on its diagonal.Similarly, we use the cosine similarity and cross entropy-based approaches to conduct contrastive learning.
@@
The hyperparameters used for all three variants of KnowCoL approach, namely KnowCoL-L, KnowCoL-H, KnowCoL-bigG, are listed as follows:
@@
We adopt the OpenCLIP ViT-L/14 image encoder as our backbone and train for 15 epochs with a batch size of 4096. Optimization is performed using AdamW (learning rate $\beta_{1}=1.0$, $\beta_{1}=1.0$, $\beta_{2}=1.0$, weight decay $1\times 10^{-4}$) together with a cosine-annealing learning-rate schedule. We set the embedding dimension $d_{e}$ to 768 and the temperature of contrastive learning $\tau$ to 0.07. In our knowledge-graph module, each entity is associated with up to 50 triplets, and we randomly sample 25 negative triplets per entity. The model was trained using 4 NVIDIA H100 GPUs, each with 80GB of memory, over a total duration of approximately 22 hours. The training was performed using mixed precision (FP16).
@@ Methods
Table 8: Example KnowCoL’s predictions on five test images. For each query, the model’s top-3 (according to the cosine similarity) Wikidata items predictions are shown (columns 1–3), with green borders indicating correct matches and red borders indicating incorrect ones; the ground-truth item is in the rightmost column.
To complement our quantitative evaluation, we present qualitative examples of KnowCoL’s predictions on the OVEN test set in Table 8. These cases illustrate the model’s ability to correctly recognize fine-grained entities under zero-shot settings.

### v2

@@ Acknowledgments
The following sections present detailed descriptions of the datasets and training procedures for each experimental scenario. We outline the experimental settings for each variant of our approach, including different fusion strategies for image and text embeddings, the score functions employed by various knowledge graph embedding (KGE) methods, and key hyperparameters. Additionally, we introduce our approach for extracting efficient subgraphs from Wikidata. Finally, we provide representative examples of multi-modal prior knowledge utilized in our KnowCoL approach for a clearer overview. The pseudo code for KnowCoL algorithms is listed in Algorithm 2 and  3.
@@
OVEN dataset consists of 14 existing datasets, grounding their labels to Wikidata IDs. These 14 datasets are originally created for image recognition/retrieval and visual question answering. They are ImageNet21k-P (Ridnik et al. 2021), iNaturalist2017 (Horn et al. 2018), Cars196 (Krause et al. 2013), SUN397 (Xiao et al. 2010), Food101 (Bossard, Guillaumin, and Gool 2014), Sports100, Aircraft (Maji et al. 2013), Oxford Flower (Nilsback and Zisserman 2008), Google Landmarks v2 (Weyand et al. 2020), VQA v2 (Goyal et al. 2019), Visual7W (Zhu et al. 2016), Visual Genome (Krishna et al. 2017), OK-VQA (Marino et al. 2019), and Text-VQA (Singh et al. 2019).
@@
In this section, we elaborate on the different fusion functions in detail, namely addition, multi-layer perception (MLP), transformer encoder (TE), and transformer encoder with image patch embeddings (TE+IP).
@@
The default fuser we used is a simple addition operation, as it shows the best generalization ability, especially for unseen entities. The fused embedding is defined as follows:
@@
The variant transformer encoder (TE) fuses the image embedding and the text embedding with a transformer encoder on top. Both embeddings are fed into the encoder, allowing them to attend to each other through self-attention. The first output embedding from the encoder is used as a global representation, capturing the fused multi-modal information.
@@
Compared to the previous approach, the local image patch embeddings from the image encoder are also fed into the transformer encoder. Similarly, the first output embedding from the encoder is used as a global representation, capturing the fused information.
@@
Here, $\phi(\cdot)$ and $\psi(\cdot)$ are functions that map nodes and relation types to corresponding embeddings.
@@
TransH is an extension for TransE, which defines relation-specific hyperplane instead of direct translation in the entity space. It can be realized by altering the score function with
@@
where $\mathbf{w}_{r}$ defines the normal vector of the hyperplane. We still use the cosine similarity and cross-entropy-based approaches to conduct contrastive learning.
@@
where $\mathrm{diag}(\mathbf{r})\in\mathbb{R}^{d\times d}$ is the diagonal matrix with $\psi(r)$ on its diagonal.Similarly, we use the cosine similarity and cross-entropy-based approaches to conduct contrastive learning.
@@
The hyperparameters used for all three variants of the KnowCoL approach, namely KnowCoL-L, KnowCoL-H, and KnowCoL-bigG, are listed as follows:
@@
We adopt the OpenCLIP ViT-L/14 image encoder as our backbone and train for 15 epochs with a batch size of 4096. Optimization is performed using AdamW (learning rate 0.0001, $\beta_{1}=1.0$, $\beta_{2}=1.0$, weight decay $1\times 10^{-4}$) together with a cosine-annealing learning-rate schedule. We set the embedding dimension $d_{e}$ to 768 and the temperature of contrastive learning $\tau$ to 0.07. In our knowledge-graph module, each entity is associated with up to 50 triplets, and we randomly sample 25 negative triplets per entity. The model was trained using 4 NVIDIA H100 GPUs, each with 80GB of memory, over a total duration of approximately 22 hours. The training was performed using mixed precision (FP16).
@@ Methods
Table 8: Example KnowCoL’s predictions on five test images. For each query, the model’s top-3 (according to the cosine similarity) Wikidata items predictions are shown, with green borders indicating correct matches and red borders indicating incorrect ones; the ground-truth item is in the rightmost column.
To complement our quantitative evaluation, we present qualitative examples of KnowCoL’s predictions on the OVEN test set in Table 8. These cases illustrate the model’s ability to recognize fine-grained entities under zero-shot settings correctly.